OpenStudy (kropot72):

Why do some of the lower case Greek characters not form when the Equation function is used? For example the lower case pi appears and remains as 'pi' whereas the upper case Pi forms the Greek character. I just saw an answer in the Mathematics Group where the answerer appeared to have the same problem as I have. Testing shows that about half of the lower case characters form correctly and all of the upper case characters form properly.

OpenStudy (lgbasallote):

hmmm i tried doing the uppercase-lower case..seems both things work IF it's witten as \[\ then the pi enclosed by \] i.e. \[\Pi\] and \[\pi\] then the words...\[Pi\] and \[pi\]

OpenStudy (lgbasallote):

what i said that transformed to latex....."seems both things work IF it's witten as...

OpenStudy (lgbasallote):

\[\

OpenStudy (lgbasallote):

then pi...then enclosed by \] (had to separate it coz it kept transforming)

OpenStudy (kropot72):

Thank you for you quick reply. I'll try your suggestion and post back.

OpenStudy (kropot72):

This is the result of trying your suggestion on pi and then directly selecting the whole set of lower case characters. \[\[\pi\]\alpha \beta \gamma \delta \epsilon \zeta \eta \theta \iota \kappa \lambda mu nu xi pi \rho \sigma \tau \upsilon \phi \chi \psi \omega\]

OpenStudy (kropot72):

Note that about half of the characters formed in the Equation preview but this was not transferred the the reply. Perhaps I should have put a space between them?

OpenStudy (lgbasallote):

regarding the pi...you did enclose it with \] right? coz the beginning thing is red...

OpenStudy (lgbasallote):

ahh they;re with pi?

OpenStudy (lgbasallote):

lemme try... \[\pi \alpha \beta \gamma \delta \epsilon \zeta \eta \theta \iota \kappa \lambda mu nu xi pi \rho \sigma \tau \upsilon \phi \chi \psi \omega\]

OpenStudy (kropot72):

Yep. I'll do it again: \[\[\pi\]\]

OpenStudy (lgbasallote):

there shouldnt be any space between pi and the \]

OpenStudy (kropot72):

Understood. I don't remember putting a space but I'll try again being extra careful: \[\[pi\]\] It happened again. I'll try the first line of lower case characters: \[\alpha \beta \gamma \delta \epsilon \zeta\]

OpenStudy (lgbasallote):

hmmm the pi is a problem...but i think i know where you're doing it wrong...are you writing it in the equations thing??

OpenStudy (kropot72):

Ohh. Yes. I'll try it in the answer box: \[\pi\] Success.

OpenStudy (lgbasallote):

:DDD

OpenStudy (kropot72):

Sorree! Now is there a way to get \[\pi\] in the Equations thing? BTW that line of Greek characters in my earlier reply were, of course, entered in the Equation thingy.

OpenStudy (lgbasallote):

ah yes. just get rid of the \[\] in the beginning and the end because equations thingy enters that automatically...so you just input \pi in the equations

OpenStudy (lgbasallote):

argh..it transformed again...get rid of \[

OpenStudy (lgbasallote):

and \]

OpenStudy (kropot72):

\[\pi\]comes up when I enter just \pi in the Equation box. Many thanks for your help. I have become a fan of yours :)

OpenStudy (lgbasallote):

<tips hat>

OpenStudy (shadowfiend):

Huh. Interesting. That should actually automatically work with no slash in the equation editor. I'll look into it.